# General geometry realized by four-scalar model and application to $f(Q)$ gravity

### v1

@@ IV.2 Spherically symmetric spacetime
Then we obtain a model that realizes the spacetime given by the metric (5) with $\nu=\nu(r)$ and $\lambda=\lambda(r)$ in the framework of $f(Q)$ gravity coupled with four scalar fields.
@@ V Summary and discussion
In this paper, we considered the reconstruction problem by using the model with four scalar fields coupled with general gravity theories in (III). The model can be a generalization of the two-scalar model (II) proposed in Nojiri:2020blr . The two-scalar model can realize any given spherically symmetric static/time-dependent spacetime in a wide class of gravity theories as shown in (20). In the four-scalar model, any given spacetime appears as a solution even if the spacetime does not have a spherical symmetry or any other symmetry. The two-scalar model includes ghosts, which make the model physically unacceptable. In the recent works Nojiri:2023dvf ; Nojiri:2023zlp ; Elizalde:2023rds ; Nojiri:2023ztz , however, it has been shown that the ghosts can be eliminated by constraints (8) given by the Lagrange multiplier fields (7). Due to the constraints, the scalar fields become non-dynamical and they do not propagate. Therefore there is not any sound, which is usually generated by the density fluctuation or oscillation of the fluid. Although the scalar fields play the role of the effective fluid, the behavior is rather different from the standard fluid, and the effective fluid is frozen. This situation is similar to the mimetic theory in Chamseddine:2013kea , where effective dark matter appears from the mimetic constraint but the effective dark matter has properties different from the standard dark matter. The mimetic dark matter is non-dynamical and even under gravitational force, it does not collapse. Even in the four-scalar model in this paper, the ghosts are eliminated by the constraints (15). Therefore, the model in this paper could be the most general extension of the mimetic theory from the effective dark matter to the effective non-dynamical arbitrary fluid including exotic fluid like a phantom.
@@
We investigated the application of the four-scalar model to $f(Q)$ gravity theory in the coincident gauge. It is well-known Banerjee:2021mqk ; Lin:2021uqa ; Wang:2021zaz that in the case that $f(Q)$ is a non-linear function of $Q$, spherically symmetric spacetime cannot be realized except in the case that $Q$ is a constant. The situation does not change if we use the two-scalar model, as we have shown in (35) but in the framework of the four-scalar model in this paper, spherically symmetric spacetime can be realized.

### v2

@@ IV.2 Spherically symmetric spacetime
Then we obtain a model that realizes the spacetime given by the metric (5) with $\nu=\nu(r)$ and $\lambda=\lambda(r)$ in the framework of $f(Q)$ gravity coupled with four scalar fields. This could tell that the diffeomorphism invariance is broken in the coincident gauge but the four scalar fields compensate for the invariance. This does not mean that the diffeomorphism invariance is recovered but in the different coordinate choices, there are corresponding different theories of four scalar fields.
@@ V Summary and discussion
In this paper, we considered the reconstruction problem by using the model with four scalar fields coupled with general gravity theories in (III). The model can be a generalization of the two-scalar model (II) proposed in Nojiri:2020blr . The two-scalar model can realize any given spherically symmetric static/time-dependent spacetime in a wide class of gravity theories as shown in (20). In the four-scalar model, any given spacetime appears as a solution even if the spacetime does not have a spherical symmetry or any other symmetry. The two-scalar model includes ghosts, which make the model physically unacceptable. In the recent works Nojiri:2023dvf ; Nojiri:2023zlp ; Elizalde:2023rds ; Nojiri:2023ztz , however, it has been shown that the ghosts can be eliminated by constraints (8) given by the Lagrange multiplier fields (7). Due to the constraints, the scalar fields become non-dynamical and they do not propagate. Therefore there is not any sound, which is usually generated by the density fluctuation or oscillation of the fluid. Although the scalar fields play the role of the effective fluid, the behaviour is rather different from the standard fluid, and the effective fluid is frozen. This situation is similar to the mimetic theory in Chamseddine:2013kea , where effective dark matter appears from the mimetic constraint but the effective dark matter has properties different from the standard dark matter. The mimetic dark matter is non-dynamical and even under gravitational force, it does not collapse. Even in the four-scalar model in this paper, the ghosts are eliminated by the constraints (15). Therefore, the model in this paper could be the most general extension of the mimetic theory from the effective dark matter to the effective non-dynamical arbitrary fluid including exotic fluid like a phantom.
@@
We investigated the application of the four-scalar model to $f(Q)$ gravity theory in the coincident gauge. It is well-known Banerjee:2021mqk ; Lin:2021uqa ; Wang:2021zaz ; Hohmann:2019fvf ; DAmbrosio:2021zpm ; Bahamonde:2022zgj ; Zhao:2021zab that in the case that $f(Q)$ is a non-linear function of $Q$, spherically symmetric spacetime cannot be realized in the polar coordinate sysytem except in the case that $Q$ is a constant. The situation does not change if we use the two-scalar model, as we have shown in (39) but in the framework of the four-scalar model in this paper, spherically symmetric spacetime can be realized.
We will now explain why we need four scalar fields to reconstruct the model that realizes a general metric. We will consider a more general scalar-tensor theory with the following action
$S_{\phi}=\int d^{4}x\sqrt{-g}\left\{\frac{1}{2}\sum_{a,b}A_{(ab)}\left(\phi^{(c)}\right)g^{\mu\nu}\partial_{\mu}\phi^{(a)}\partial_{\nu}\phi^{(b)}-V\left(\phi^{(c)}\right)\right\}\,.$ (45)
Here we include $N$ scalar fields, $\phi^{(a)}$ $\left(a=1,2,\cdots,N\right)$. Note that there is an ambiguity in the redefinition of the scalar fields $\phi^{(a)}\to{\tilde{\phi}}^{(a)}\left(\phi^{(b)}\right)$, which changes $A_{(ab)}$ by $A_{(ab)}\left(\phi^{(c)}\right)\to{\tilde{A}}_{(ab)}\left({\tilde{\phi}}^{(c)}\right)\equiv\sum_{d,f}\frac{\partial\phi^{(d)}}{\partial{\tilde{\phi}}^{(a)}}{\partial\phi^{(f)}}{\partial{\tilde{\phi}}^{(b)}}A_{(df)}\left(\phi^{(h)}\left({\tilde{\phi}}^{(c)}\right)\right)$. This is an analogue of a coordinate transformation.
When $N=1$, through a change of variables, we can set $A_{(ab)}$ to be ±$\pm 1$. After this adjustment, only one function remains, that is, the potential $V(\phi)$. This is a key factor in how the single-scalar model can reconstruct a model that realizes the general FLRW universe Nojiri:2005pu ; Capozziello:2005tf .
In spherically symmetric spacetime, the metric contains three functions: $g_{tt}$, $g_{tr}=g_{rt}$, and $g_{rr}$. Frequently, we can eliminate $g_{tr}$ through a coordinate transformation, as shown in (5). Therefore, at least two scalar fields are necessary to realize the spherically symmetric spacetime in the framework of the scalar-tensor theory Nojiri:2020blr . In the two-scalar theory, there are three components of $A_{(ab)}\left(\phi^{(c)}\right)$ and a potential $V\left(\phi^{(c)}\right)$. Due to the ambiguity of the field redefinition with two functional degrees of freedom, two functional degrees of freedom remain among the four functional degrees of $A_{(ab)}\left(\phi^{(c)}\right)$ and $V\left(\phi^{(c)}\right)$.
In a general four-dimensional spacetime, the metric possesses ten components, with an additional four redundancies stemming from coordinate transformations. This indicates that, to encompass the general spacetime within the framework of scalar-tensor theory outlined in (45), we require a minimum of four scalar fields. These fields give rise to ten functions $A_{(ab)}\left(\phi^{(c)}\right)$ and a potential $V\left(\phi^{(c)}\right)$, accompanied by four redundancies resulting from the redefinition of scalar fields. Therefore, the scalar-tensor theory (45) possesses one additional functional degree of freedom. Consequently, selecting $V=0$ yields a non-linear $\sigma$ model, with its target space being a four-dimensional manifold.